OpenStudy (anonymous):

Rachel went hiking for one hundred fifty minutes. Express this time in hour and minutes.

OpenStudy (anonymous):

how many minutes are in an hour?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

you divide 150 minutes by 60 because there are 60minutes in an hour. From there you can get your answer

OpenStudy (anonymous):

Write back once you have solved it :)

OpenStudy (anonymous):

I want to help him

OpenStudy (anonymous):

@nataliedhatshortiee did you get your answer yet?

OpenStudy (anonymous):

2hour 5min @AwesomeB

OpenStudy (anonymous):

I got 2 hours and 30 minutes

OpenStudy (jdoe0001):

1 hour = 60mins 2hours = 60 + 60 = 120 mins + 5mins 120 + 5 = 125mins

OpenStudy (anonymous):

just add the stuff up